# Listen, Look, Drive: Coupling Audio Instructions for User-aware VLA-based Autonomous Driving

### v2

@@ Abstract
Vision Language Action (VLA) models promise an open-vocabulary interface that can translate perceptual ambiguity into semantically grounded driving decisions, yet they still treat language as a static prior fixed at inference time. As a result, the model must infer continuously shifting objectives from pixels alone, yielding delayed or overly conservative maneuvers. We argue that effective VLAs for autonomous driving need an online channel in which users can influence driving with specific intentions. To this end, we present EchoVLA, a user-aware VLA that couples camera streams with in situ audio instructions. We augment the nuScenes dataset with temporally aligned, intent-specific speech commands generated by converting ego-motion descriptions into synthetic audios. Further, we compose emotional speech-trajectory pairs into a multimodal Chain-of-Thought (CoT) for fine-tuning a Multimodal Large Model (MLM) based on Qwen2.5-Omni. Specifically, we synthesize the audio-augmented dataset with different emotion types paired with corresponding driving behaviors, leveraging the emotional cues embedded in tone, pitch, and speech tempo to reflect varying user states, such as urgent or hesitant intentions, thus enabling our EchoVLA to interpret not only the semantic content but also the emotional context of audio commands for more nuanced and emotionally adaptive driving behavior. In open-loop benchmarks, our approach reduces the average L2 error by $59.4\%$ and the collision rate by $74.4\%$ compared to the baseline of vision-only perception. More experiments on nuScenes dataset validate that EchoVLA not only steers the trajectory through audio instructions, but also modulates driving behavior in response to the emotions detected in the user’s speech.
@@ III-A Multimodal CoT
Based on the ratios between the synthetic and original arousal values, we modulate the corresponding trajectories with a faster and slower speed profile without changing the original driving directions. Given an input trajectory $\mathbf{P}=\{\mathbf{p}_{i}\in\mathbb{R}^{2}\}_{i=0}^{N-1}$ sampled uniformly over time $t\in[0,T]$ and a target emotion $e\in\{\text{urgent},\text{hesitant}\}$, we define the modulation procedure to synthesize a new trajectory $\tilde{\mathbf{P}}=(\tilde{p}_{x},\tilde{p}_{y})$ according to an emotion-conditioned speed profile as follows.
@@ IV-C Evaluation Methods
To verify our proposed method, we compared our EchoVLA with visual perception-only VLA on the nuScenes open-loop L2 error and collision rate. The baseline visual perception-only VLA is based on a larger VLM Qwen2-VL-7B compared to our 3B scale model. In Qwen2-VL, Multimodal Rotary Position Embedding (M-RoPE) is applied to process multimodal input by decomposing rotary embedding into temporal, height, and width components, which equips Qwen2-VL with powerful multimodal data handling capabilities [27]. The baseline visual perception-only VLA is also fine-tuned on the nuScenes dataset.
@@ V Conclusion
We introduce EchoVLA, which equips VLA with an online audio channel, allowing users to inject time-varying intent without hand-engineered reward reshaping or additional advanced sensors. By converting ego-motion descriptions into synchronized speech commands and fine-tuning a Qwen2.5-Omni backbone, EchoVLA learns to align acoustic intent with visual context and regresses trajectories in a CoT reasoning way. The open-loop evaluation in nuScenes confirms that audio enhancement is complementary to, rather than competing with, visual perception.
However, current audio command generation relies on synthesized speech derived from recorded ego trajectories, which does not fully capture the variability of real human speech. In practice, user utterances are often noisier, span multiple languages, and can be ambiguous in meaning. In addition, the model processes the entire audio clip before generating a response, increasing latency for longer commands—streaming token generation has not yet been implemented. Our system assumes clean audio input and has not been evaluated under realistic acoustic conditions such as heavy traffic noise, wind, or background music playing inside the cabin.
Looking ahead, we plan to address these limitations by collecting a large-scale corpus of real in-car voice commands that reflect diverse languages, accents, and paraphrased expressions. Finally, we also intend to enhance robustness through an adversarial noise-training phase and by incorporating an on-board denoising front-end to handle challenging acoustic environments.

### v3

@@ Abstract
Vision Language Action (VLA) models promise an open-vocabulary interface that can translate perceptual ambiguity into semantically grounded driving decisions, yet they typically treat language as a static prior fixed at inference time. As a result, the model must infer continuously shifting objectives from pixels alone, yielding delayed or overly conservative maneuvers. We argue that effective VLAs for autonomous driving need an online channel in which users can influence driving with specific intentions. To this end, we present EchoVLA, a user-aware VLA that couples camera streams with in situ audio instructions. We augment the nuScenes dataset with temporally aligned, intent-specific speech commands generated by converting ego-motion descriptions into synthetic audios. Further, we compose emotional speech-trajectory pairs into a multimodal Chain-of-Thought (CoT) for fine-tuning a Multimodal Large Model (MLM) based on Qwen2.5-Omni. Specifically, we synthesize the audio-augmented dataset with different emotion states paired with corresponding driving behaviors, leveraging the emotional cues embedded in tone, pitch, and speech tempo to reflect varying user states, such as urgent or hesitant intentions, thus enabling our EchoVLA to interpret not only the semantic content but also the emotional context of audio commands for more nuanced and emotionally adaptive driving behavior. In open-loop benchmarks, our approach reduces the average L2 error by $59.4\%$ and the collision rate by $74.4\%$ compared to the baseline of vision-only perception. More experiments on nuScenes dataset validate that EchoVLA not only steers the trajectory through audio instructions, but also modulates driving behavior in response to the emotions detected in the user’s speech.
@@ III-A Multimodal CoT
Based on the ratios between the synthetic and original arousal values, we modulate the corresponding trajectories with a faster or slower speed profile without changing the original driving directions. Given an input trajectory $\mathbf{P}=\{\mathbf{p}_{i}\in\mathbb{R}^{2}\}_{i=0}^{N-1}$ sampled uniformly over time $t\in[0,T]$ and a target emotion $e\in\{\text{urgent},\text{hesitant}\}$, we define the modulation procedure to synthesize a new trajectory $\tilde{\mathbf{P}}=(\tilde{p}_{x},\tilde{p}_{y})$ according to an emotion-conditioned speed profile as follows.
@@ IV-C Evaluation Methods
To validate our proposed method, we compared our EchoVLA with visual perception-only VLA on the nuScenes open-loop L2 error and collision rate. The baseline visual perception-only VLA is based on a larger VLM Qwen2-VL-7B compared to our 3B scale model. In Qwen2-VL, Multimodal Rotary Position Embedding (M-RoPE) is applied to process multimodal input by decomposing rotary embedding into temporal, height, and width components, which equips Qwen2-VL with robust multimodal data handling capabilities [27]. The baseline visual perception-only VLA is also fine-tuned on the nuScenes dataset.
@@ V Conclusion
We introduce EchoVLA, which augments VLA with an online audio channel, allowing users to inject time-varying intent without hand-engineered reward reshaping or additional advanced sensors. By converting ego-motion descriptions into synchronized speech commands and fine-tuning a Qwen2.5-Omni backbone, EchoVLA learns to align acoustic intent with visual context and regresses trajectories in a CoT reasoning way. The open-loop evaluation in nuScenes confirms that audio enhancement is complementary to, rather than competing with, visual perception.
However, current audio command generation relies on synthesized speech derived from recorded ego trajectories, which does not fully capture the variability of natural human speech. In practice, user utterances are often noisier, span multiple languages, and can be ambiguous in meaning. Additionally, the model processes the entire audio clip before generating a response, increasing latency for longer commands—streaming token generation has not yet been implemented. Our system assumes clean audio input and has not been evaluated under realistic acoustic conditions such as heavy traffic noise, wind, or background music playing inside the cabin.
Moving forward, we plan to address these limitations by curating a large-scale corpus of real in-car voice commands that reflect diverse languages, accents, and paraphrased expressions. We also intend to enhance robustness through an adversarial noise-training phase and by incorporating an on-board denoising front-end to handle challenging acoustic environments.